My answers are incorrect. need helpThe structure shown below is supported by a pin and roller on an inclined surface that is parallel to the line connecting joi
Answer #1

soll! The distance beth, pin A and rollen G along the inclined plane by using Pythagoras theorem. . AG² = AH? + Gim? AG = VAHof Moments about G; considen the equiebrium EMG=0 (Px18)-(Px/8)-(Px36)-(PX54)+(Ax * 12)+(Ay x54)=0 >(480x18) --C 480 X16)-(48Ages-3381.843 lb Axes 338.843 46 6). from ego, 12 Ax+ 54 Ay 43200 =0 (12 X-338-843 ) + Su Ay. = 43200 Aya 4+266.116 54 . Ay =FITS. 11181.418 12 FIJ = 931-79 2b (Tension) considen equilibrium of forces in tndinn: . Efx=0 Foot FCI cos 33.69* + Fpj + A